# Self-Supervised Uncalibrated Multi-View Video Anonymization in the Operating Room

### v1

@@ 1 Introduction
We evaluate our approach on both the 4D-OR dataset [ozsoy20224d] of simulated surgeries and our collected dataset of real surgeries in a more complex scene. For each person, we annotate (1) the whole-body bounding box, (2) the “hard case” flag (over 67% occlusion), and (3) three keypoints (eyes and chin) if they are visible. We also introduce two new metrics to quantify robustness against challenging scenes: (1) hard case recall, which calculates the recall rate of the hard cases; (2) multi-view recall, which measures the percentage of subjects anonymized across all camera views in which they appear. Experimental results show that our approach outperforms existing methods, specifically by achieving a recall rate over 97% on both datasets. Compared to state-of-the-art approaches, ours reduces 13,500 missed detections in a one-hour real surgical video, which saves significant manual reviewing time.
We summarize our contributions as follows: (1) We address video anonymization in the OR without using any annotations. (2) We propose a two-stage anonymization pipeline that detects whole bodies and estimates whole-body keypoints, which achieves state-of-the-art performance and supports anonymization of any desired body region. (3) We propose to combine tracking and multi-view person association for more robust whole-body detection using temporal and multi-view context. (4) We solve domain adaptation and real-time detection by using spatial-temporally augmented detections as pseudo labels for model fine-tuning. (5) We propose hard case recall and multi-view recall as two new metrics, and evaluate different approaches on both actor-simulated and real surgical videos at three anonymization levels: whole-body, face, and eye.
@@ 3.3 Self-supervised multi-view person association
P R AP@.5 $\text{R}_{\mathrm{hard}}$ P R AP@.3 $\text{R}_{\mathrm{hard}}$ MvR P R AP@.3 $\text{R}_{\mathrm{hard}}$ MvR
@@
P R AP@.5 $\text{R}_{\mathrm{hard}}$ P R AP@.3 $\text{R}_{\mathrm{hard}}$ MvR P R AP@.3 $\text{R}_{\mathrm{hard}}$ MvR
@@ 3.5 Whole-body pose estimation
P R AP@.5 $\text{R}_{\mathrm{hard}}$ P R AP@.3 $\text{R}_{\mathrm{hard}}$ MvR P R AP@.3 $\text{R}_{\mathrm{hard}}$ MvR
@@ 4.1 Datasets and evaluation metrics
We evaluate our approach using five metrics: precision (P), recall (R), average precision (AP), hard case recall ($\text{R}_{\mathrm{hard}}$), and multi-view recall (MvR). Apart from standard AP, we additionally report P, R, and $\text{R}_{\mathrm{hard}}$ when the score threshold is set to 0.1, because in the context of video anonymization, R is more important than P in order not to miss a person, and $\text{R}_{\mathrm{hard}}$ explicitly evaluates models’ capability to detect heavily-occluded cases. To further showcase the unique challenges of multi-camera setups, we introduce MvR, a stricter metric that measures the percentage of subjects successfully anonymized across all camera views in which they appear. A subject is only considered correctly anonymized if they are detected in every single view of a given multi-view set.
@@
P R AP@.5 $\text{R}_{\mathrm{hard}}$ P R AP@.3 $\text{R}_{\mathrm{hard}}$ MvR P R AP@.3 $\text{R}_{\mathrm{hard}}$ MvR
@@ 4.2 Results
P R $\text{R}_{\mathrm{hard}}$ MvR P R $\text{R}_{\mathrm{hard}}$ MvR

### v2

@@ 1 Introduction
We evaluate our approach on both the 4D-OR dataset [ozsoy20224d] of simulated surgeries and our collected dataset of real surgeries in a more complex scene. For each person, we annotate (1) the whole-body bounding box, (2) the “hard case” flag (over 67% occlusion), and (3) three keypoints (eyes and chin) if they are visible. We also apply two strict metrics to quantify robustness against challenging scenes: (1) hard case recall, which calculates the recall rate of the hard cases; (2) holistic recall, which measures the percentage of subjects anonymized across all camera views in which they appear [bastian2023disguisor]. Experimental results show that our approach outperforms existing methods, specifically by achieving a recall rate over 97% on both datasets. Compared to state-of-the-art approaches, ours reduces 13,500 missed detections in a one-hour real surgical video, which saves significant manual reviewing time.
We summarize our contributions as follows: (1) We address video anonymization in the OR without using any annotations. (2) We propose a two-stage anonymization pipeline that detects whole bodies and estimates whole-body keypoints, which achieves state-of-the-art performance and supports anonymization of any desired body region. (3) We propose to combine tracking and multi-view person association for more robust whole-body detection using temporal and multi-view context. (4) We solve domain adaptation and real-time detection by using spatial-temporally augmented detections as pseudo labels for model fine-tuning. (5) We use strict metrics including hard case recall and holistic recall, and evaluate different approaches on both actor-simulated and real surgical videos at three anonymization levels: whole-body, face, and eye.
@@ 3.3 Self-supervised multi-view person association
P R AP@.5 $\text{R}_{\mathrm{hard}}$ P R AP@.3 $\text{R}_{\mathrm{hard}}$ HoR P R AP@.3 $\text{R}_{\mathrm{hard}}$ HoR
@@
P R AP@.5 $\text{R}_{\mathrm{hard}}$ P R AP@.3 $\text{R}_{\mathrm{hard}}$ HoR P R AP@.3 $\text{R}_{\mathrm{hard}}$ HoR
@@ 3.5 Whole-body pose estimation
P R AP@.5 $\text{R}_{\mathrm{hard}}$ P R AP@.3 $\text{R}_{\mathrm{hard}}$ HoR P R AP@.3 $\text{R}_{\mathrm{hard}}$ HoR
@@ 4.1 Datasets and evaluation metrics
We evaluate our approach using five metrics: precision (P), recall (R), average precision (AP), hard case recall ($\text{R}_{\mathrm{hard}}$), and holistic recall (HoR). Apart from standard AP, we additionally report P, R, and $\text{R}_{\mathrm{hard}}$ when the score threshold is set to 0.1, because in the context of video anonymization, R is more important than P in order not to miss a person, and $\text{R}_{\mathrm{hard}}$ explicitly evaluates models’ capability to detect heavily-occluded cases. To further showcase the unique challenges of multi-camera setups, we apply HoR [bastian2023disguisor], where a subject is only considered correctly anonymized if they are detected in every single view of a given multi-view set if visible.
@@
P R AP@.5 $\text{R}_{\mathrm{hard}}$ P R AP@.3 $\text{R}_{\mathrm{hard}}$ HoR P R AP@.3 $\text{R}_{\mathrm{hard}}$ HoR
@@ 4.2 Results
P R $\text{R}_{\mathrm{hard}}$ HoR P R $\text{R}_{\mathrm{hard}}$ HoR